OpenStudy (anonymous):

Evaluate this limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2} \frac{ 4-x^2 }{ 2-x }\]

OpenStudy (anonymous):

Hint: It's an indeterminate form.

OpenStudy (anonymous):

i know it's 0/0 but i don't know what to do next

OpenStudy (anonymous):

You also have difference of squares!

OpenStudy (anonymous):

\[(a^2-b^2)=(a-b)(a+b)\]

OpenStudy (anonymous):

We calculate the limit substituting x by 2 lim (4 - x^2 )/(2-x) = (4-4)/(2-2) = 0/0 We could write the numerator as a difference of squares: 4 - x^2 = (2 - x)(2 + x) lim (4 - x^2)/(2-x) = lim (2 - x)(2 + x)/(2-x) We'll simplify and we'll get: lim (2+x) = 4

OpenStudy (raden):

anyway u may use derivative derivative of 4-x^2 is -2x and derivative of 2-x is -1 so, just plug in x=2 to (-2x)/-1 = 4

OpenStudy (anonymous):